Presentation is loading. Please wait.

Presentation is loading. Please wait.

4th July Two physics fill during the day Smooth data taking LHC plans changed: 90m run postponed to tomorrow (24h) OT FE still excluded, to be looked at.

Similar presentations


Presentation on theme: "4th July Two physics fill during the day Smooth data taking LHC plans changed: 90m run postponed to tomorrow (24h) OT FE still excluded, to be looked at."— Presentation transcript:

1 4th July Two physics fill during the day Smooth data taking LHC plans changed: 90m run postponed to tomorrow (24h) OT FE still excluded, to be looked at in more details today during the cryo intervention TPU TELL1 throttling

2 LHC LHC LAST 24H AND PLANS (July 5) - Wednesday morning - afternoon: Stable Beams declared for fill 2806 with 1374 bunches at 09:30, dumped at 18:50. - Wednesday evening: Two attempts to fill for physics in the evening lost during injection due to injection failures with bunches hitting straight the injection absorber. Bunches arrived 50us too early due to SPS RF being on local clock. It is not understood why this was not caught by interlock, to be investigated. - Wednesday evening: Oil leak discovered on surface screw cryo compressor at point 2, intervention scheduled for Thursday morning for 4h. - Wednesday night: Stable Beams declared with 1374 bunches at midnight. Oil leak on screw compressor worsened, and finally compressor lost at 03:40. System reconfigured to survive with only one screw compressor, cryo conditions not lost, down in shadow while beams in physics. Plan - Attempt repairing screw compressor at point 2 in the shadow of physics fill, if the reweld disturbes neighbouring equipment could lead to loss of cryo conditions and 12h recovery. If not, repair would be finished by noon with no loss of cryo and disturbance to beam. - Keep current fill. - Refill for physics late morning-afternoon depending on success of compressor repair. 90m beta* setup starting Friday morning ~09:00 for 24h - Resume physics on Saturday morning

3 Fill 2806 Started 9:00, ended 18: colliding bunches,  =1.7, L=3.9x10 32 Record 12 pb -1, with 94% efficiency L0 raw rate: 937 kHz HLT Physics rate : 4 kHz Deferral 23%, all deferred runs processed after fill. Problems: – OT FE disabled – Some farm nodes excluded/included

4 Fill 2807 Started midnight ongoing 1262 colliding bunches,  =1.7, L=3.9x pb -1 recorded with 95% efficiency Problems: – OT FE desactivated – Some Farm nodes excluded/included – Manual reset of RICH to recover disabled HPD took some time to do


Download ppt "4th July Two physics fill during the day Smooth data taking LHC plans changed: 90m run postponed to tomorrow (24h) OT FE still excluded, to be looked at."

Similar presentations


Ads by Google